given a,b,c are three vectors, then [a,b,c] is not equal to?
a) [a,c,b]
b)[c,a,b]
c) -[a,c,b]

acc to me (a) is not equal since it is not cylic permutated.. but in the answer key (c) is the right option.. howcome..?

Grade:12th pass

2 Answers

Arun Kumar IIT Delhi
9 years ago
[a,b,c]=a*(b*c)
[a,c,b]=a*(c*b)=-(a*(b*c))=-[a,b,c]
RAJAT CHHABRA
9 years ago
what is your answer. i didnt get it